Hi Thanks for stopping by!!! Here is my question: Perform the indicated operation

7/5 . 8/3
These problems are supposed to look like fractions, but I was not able to make them look like fractions I hope this looks right? Thanks for your help!

Answer by stanbon(75887)

7/5 . 8/3
----
If you are multiplying the fractions
you get 56/15
---
If you are dividing the fractions
you get (7/5)/(8/3) = (7/5)*(3/8) = 21/40
================================================
Cheers,
Stan H.